OpenStudy (anonymous):

Please help me solve triangle ABC (sine/cosine theorem) if there is: b + c = 20 ; a = 5*sqrt(2) , γ = 135°

OpenStudy (anonymous):

if \[\alpha + \beta + \gamma = 180 ° \] how to find b and c through sine theorem? \[\frac{ a }{ \sin \alpha } = \frac{ b }{ \sin \beta } = \frac{ c }{ \sin \gamma }\] There is no alpha or beta angles.